OpenStudy (anonymous):

Evaluate. 7!/5!2! 1 21 42

OpenStudy (anonymous):

7!/5! you can cancel out the 5, 4, 3, 2, 1, so you get 7*6, then divide by 2!=2*1=2, so 7*6/2=21